# Amortized Variational Transdimensional Inference

### v2

@@ Abstract
The expressiveness of flow-based models combined with stochastic variational inference (SVI) has expanded the application of optimization-based Bayesian inference to highly complex problems. However, despite the importance of multi-model Bayesian inference, defined over a transdimensional joint model and parameter space, flow-based SVI has been limited to problems defined over a fixed-dimensional parameter space. We introduce CoSMIC  normalizing flows (COntextually-Specified Masking for Identity-mapped Components), an extension to neural autoregressive conditional normalizing flow architectures that enables use of a single amortized variational density for inference over a transdimensional (multi-model) conditional target distribution. We propose a combined stochastic variational transdimensional inference (VTI) approach to training CoSMIC  flows using ideas from Bayesian optimization and Monte Carlo gradient estimation. Numerical experiments show the performance of VTI on challenging problems that scale to high-cardinality model spaces.
@@ 1 Introduction
Amortized variational inference [Dayan et al., 1995] has seen a surge in interest since the introduction of normalizing flows [Rezende and Mohamed, 2015]. Amortized densities can be used for a variety of downstream tasks, such as importance sampling [Rezende and Mohamed, 2015], simulation-based inference [Papamakarios et al., 2019, Zammit-Mangion et al., 2024], adaptive Markov chain Monte Carlo (MCMC) [Gabrié et al., 2022], and generative modeling [Kingma and Welling, 2013]. While many existing approaches only consider continuous supports, there is a growing interest in applications where the support is either discrete or discretely indexed [Diluvi et al., 2024]. One such application concerns a target transdimensional probability distribution $\pi$ with support $\pazocal{X}=\smash{\mathop{\textstyle\bigcup}\nolimits_{m\in\pazocal{M}}(\{m\}\times\Theta_{m})}$, where $\pazocal{M}$ is a finite discrete index set, $\Theta_{m}\subseteq\mathbb{R}^{d_{m}}$, and the dimension $d_{m}$ of $\Theta_{m}$ may vary with $m$. Hence $\pazocal{X}$ is a transdimensional space [Green, 1995, Sisson, 2005, Fan et al., 2026]. Such spaces arise in Bayesian model inference, where $\Theta_{m}$ correspond to model parameters, and $m\in\pazocal{M}$ is a model index. Discrete indices parameterize many practical inference problems, including variable selection [Fan et al., 2026], mixtures-of-regressions, learning directed acyclic graphs (DAGs) from data [Thompson et al., 2025], phylogenetic tree topology search Everitt et al. [2020], mixture-component inference Das and Bhattacharya [2014], geoscientific inversion Sambridge et al. [2013], and change-point models Green [1995]. This article is concerned with estimating the target distribution $\pi$ with associated density function $\pi(m,\bm{\theta}_{m})$, $\bm{\theta}_{m}\in\Theta_{m}$, whose dimension depends on $m$. For simplicity we refer to $\pi(m,\bm{\theta}_{m})$ and related functions as density functions, even though they are not continuous. Typically, this density is only available in a conditional unnormalized form, $\eta(\bm{\theta}_{m}\,|\,m)=\pazocal{Z}_{m}\pi(\bm{\theta}_{m}\,|\,m)$, where $\pazocal{Z}_{m}=\smash{\int_{\Theta_{m}}\eta(\bm{\theta}_{m}\,|\,m)d\bm{\theta}_{m}}$. The factorization $\eta(m,\bm{\theta}_{m})=\eta(\bm{\theta}_{m}\,|\,m)\pi(m)$ implies there is a discrete target probability mass function over models, $\pi(m)=\pazocal{Z}_{m}\pazocal{Z}^{-1}$, where $\pazocal{Z}=\smash{\sum_{m\in\pazocal{M}}\pazocal{Z}_{m}}$. Estimation of $\eta(m,\bm{\theta}_{m})$ then becomes estimation of both $\eta(\bm{\theta}_{m}\mid m)$ and $\pi(m)$.
@@
Contributions: We introduce CoSMIC  (COntextually-Specified Masking for Identity-mapped Components) flows, a widely applicable and simple modification to conditional neural flow architectures (Section 2). CoSMIC  flows fundamentally expand the use cases for normalizing flows to encompass amortized variational inference applications, so that a single amortized variational density can be used for variational inference over a transdimensional (multi-model) target distribution. In effect, this extends the reparameterization trick exploited by IAF-based VI to the transdimensional setting. In Section 3, we demonstrate the efficacy of CoSMIC  transformations within a novel variational transdimensional inference (VTI) framework with two implementations. The first builds upon principles of Bayesian optimization [Srinivas et al., 2012], and the second uses Monte Carlo gradient estimation [Mohamed et al., 2020]. We also provide a theoretical analysis of VTI approximation error bounds under a Gaussian process surrogate, and convergence guarantees for the marginal model distribution under convergent optimization steps. Finally, we demonstrate the applicability of VTI to problems with model spaces that cannot be easily enumerated within the memory limitations of current computing architectures. In particular, Section 5 explores problems in Bayesian robust variable selection [O’Hara and Sillanpää, 2009] and Bayesian causal discovery [Heckerman et al., 2006].  [^1] PyTorch CUDA code for all experiments is available at https://github.com/daviesl/avti.
@@ 5.1 Bayesian misspecified robust variable selection
Figure 3: Left: A simulation study of the robust variable selection example showing the cross entropy (NLL) between RJMCMC samples and an amortized variational transdimensional density using rational quadratic spline CoSMIC flows under a fixed number of iterations (30,000). Each cardinality was run with 10 independently sampled synthetic data sets. Right: Comparison of bivariate plots of variables $\smash{\bm{\theta}_{m}^{(1)},\bm{\theta}_{m}^{(5)}}$ obtained by RJMCMC and VTI for a single $\smash{|\pazocal{M}|=2^{7}}$ problem.
5.2 Bayesian non-linear directed acyclic graph discovery
@@ 6 Discussion
We have introduced CoSMIC  normalizing flows as a means to implement amortized variational transdimensional inference (VTI), the approximation of a target density over a transdimensional space with a single variational density. VTI is broadly applicable to a wide class of transdimensional inference problems. Although the specification of a CoSMIC flow requires augmenting all model dimensions to $d_{\mathrm{max}}$}, VTI is not sensitive to these added dimensions during training and inference due to the construction of the auxiliary variable transforms. We presented two approaches for simultaneously optimizing the variational parameters $\psi,\phi$. The Gaussian surrogate-based approach benefits from our derivation of the approximation error bounds and established convergence guarantees for the marginal models distribution under convergent optimization steps. The two approaches that use Monte Carlo gradient estimation for SGD optimization benefit from recent advances in neural architectures and neural approximation of very large model spaces. The choice of model sampler is dependent on both the cardinality of the model space and the structure of the problem. When $|\pazocal{M}|$ is small, the Gaussian process surrogate-based sampler or the categorical sampler using Monte Carlo gradients are both appropriate, although in practice it is usually safe to default to the latter approach. For high cardinality problems, we recommend a neural model sampler for approximate inference on the distribution of model weights.
@@ Acknowledgments
This research was made possible by funding support from the CSIRO Machine Learning and Artificial Intelligence Future Science Platform. The authors would also like to thank Ryan Thompson for critical advice in the construction of the non-linear DAG example, and Edwin Bonilla for advice on Monte Carlo gradient estimation methods. SAS is supported by the Australian Research Council.

### v3

# Variational Transdimensional Inference

@@ Abstract
The expressiveness of flow-based models combined with stochastic variational inference (SVI) has expanded the application of optimization-based Bayesian inference to highly complex problems. However, despite the importance of multi-model Bayesian inference for problems defined on a transdimensional joint model and parameter space, such as Bayesian structure learning and model selection, flow-based SVI has been limited to problems defined on a fixed-dimensional parameter space. We introduce CoSMIC  normalizing flows (COntextually-Specified Masking for Identity-mapped Components), an extension to neural autoregressive conditional normalizing flow architectures that enables use of a single flow-based variational density for inference over a transdimensional (multi-model) conditional target distribution. We propose a combined stochastic variational transdimensional inference (VTI) approach to training CoSMIC  flows using ideas from Bayesian optimization and Monte Carlo gradient estimation. Numerical experiments show the performance of VTI on challenging problems that scale to high-cardinality model spaces.
@@ 1 Introduction
Variational inference via stochastic optimization [Paisley et al., 2012, Hoffman et al., 2013] has surged in interest since the introduction of normalizing flows [Rezende and Mohamed, 2015]. Flow-based densities can be used for a variety of downstream tasks, such as importance sampling [Rezende and Mohamed, 2015], simulation-based inference [Papamakarios et al., 2019, Zammit-Mangion et al., 2024], adaptive Markov chain Monte Carlo (MCMC) [Gabrié et al., 2022], and generative modeling [Kingma and Welling, 2013]. While many existing approaches only consider continuous supports, there is a growing interest in applications where the support is either discrete or discretely indexed [Diluvi et al., 2024]. One such application concerns a target transdimensional probability distribution $\pi$ with support $\pazocal{X}=\smash{\mathop{\textstyle\bigcup}\nolimits_{m\in\pazocal{M}}(\{m\}\times\Theta_{m})}$, where $\pazocal{M}$ is a finite discrete index set, $\Theta_{m}\subseteq\mathbb{R}^{d_{m}}$, and the dimension $d_{m}$ of $\Theta_{m}$ may vary with $m$. Hence $\pazocal{X}$ is a transdimensional space [Green, 1995, Sisson, 2005, Fan et al., 2026]. Such spaces arise in Bayesian model inference, where $\Theta_{m}$ correspond to model parameters, and $m\in\pazocal{M}$ is a model index. Discrete indices parameterize many practical inference problems, including variable selection [Fan et al., 2026], mixtures-of-regressions, learning directed acyclic graphs (DAGs) from data [Thompson et al., 2025], phylogenetic tree topology search Everitt et al. [2020], mixture-component inference Das and Bhattacharya [2014], geoscientific inversion Sambridge et al. [2013], and change-point models Green [1995]. This article is concerned with estimating the target distribution $\pi$ with associated density function $\pi(m,\bm{\theta}_{m})$, $\bm{\theta}_{m}\in\Theta_{m}$, whose dimension depends on $m$. For simplicity we refer to $\pi(m,\bm{\theta}_{m})$ and related functions as density functions, even though they are not continuous. Typically, this density is only available in a conditional unnormalized form, $\eta(\bm{\theta}_{m}\,|\,m)=\pazocal{Z}_{m}\pi(\bm{\theta}_{m}\,|\,m)$, where $\pazocal{Z}_{m}=\smash{\int_{\Theta_{m}}\eta(\bm{\theta}_{m}\,|\,m)d\bm{\theta}_{m}}$. The factorization $\eta(m,\bm{\theta}_{m})=\eta(\bm{\theta}_{m}\,|\,m)\pi(m)$ implies there is a discrete target probability mass function over models, $\pi(m)=\pazocal{Z}_{m}\pazocal{Z}^{-1}$, where $\pazocal{Z}=\smash{\sum_{m\in\pazocal{M}}\pazocal{Z}_{m}}$. Estimation of $\eta(m,\bm{\theta}_{m})$ then becomes estimation of both $\eta(\bm{\theta}_{m}\mid m)$ and $\pi(m)$.
@@
Contributions: We introduce CoSMIC  (COntextually-Specified Masking for Identity-mapped Components) flows, a widely applicable and simple modification to conditional neural flow architectures (Section 2). CoSMIC  flows fundamentally expand the use cases for normalizing flows to encompass variational inference applications, so that a single flow-based variational density can be used for variational inference over a transdimensional (multi-model) target distribution. In effect, this extends the reparameterization trick exploited by IAF-based VI to the transdimensional setting. In Section 3, we demonstrate the efficacy of CoSMIC  transformations within a novel variational transdimensional inference (VTI) framework with two implementations. The first builds upon principles of Bayesian optimization [Srinivas et al., 2012], and the second uses Monte Carlo gradient estimation [Mohamed et al., 2020]. We also provide a theoretical analysis of VTI approximation error bounds under a Gaussian process surrogate, and convergence guarantees for the marginal model distribution under convergent optimization steps. Finally, we demonstrate the applicability of VTI to problems with model spaces that cannot be easily enumerated within the memory limitations of current computing architectures. In particular, Section 5 explores problems in Bayesian robust variable selection [O’Hara and Sillanpää, 2009] and Bayesian causal discovery [Heckerman et al., 2006].  [^1] PyTorch CUDA code for all experiments is available at https://github.com/daviesl/avti.
@@ 5.1 Bayesian misspecified robust variable selection
Figure 3: Left: A simulation study of the robust variable selection example showing the cross entropy (NLL) between RJMCMC samples and an flow-based variational transdimensional density using rational quadratic spline CoSMIC flows under a fixed number of iterations (30,000). Each cardinality was run with 10 independently sampled synthetic data sets. Right: Comparison of bivariate plots of variables $\smash{\bm{\theta}_{m}^{(1)},\bm{\theta}_{m}^{(5)}}$ obtained by RJMCMC and VTI for a single $\smash{|\pazocal{M}|=2^{7}}$ problem.
5.2 Bayesian structure learning via non-linear directed acyclic graph discovery
@@ 6 Discussion
We have introduced CoSMIC  normalizing flows as a means to implement variational transdimensional inference (VTI), the approximation of a target density over a transdimensional space with a single variational density. VTI is broadly applicable to a wide class of transdimensional inference problems. Although the specification of a CoSMIC flow requires augmenting all model dimensions to $d_{\mathrm{max}}$}, VTI is not sensitive to these added dimensions during training and inference due to the construction of the auxiliary variable transforms. We presented two approaches for simultaneously optimizing the variational parameters $\psi,\phi$. The Gaussian surrogate-based approach benefits from our derivation of the approximation error bounds and established convergence guarantees for the marginal models distribution under convergent optimization steps. The two approaches that use Monte Carlo gradient estimation for SGD optimization benefit from recent advances in neural architectures and neural approximation of very large model spaces. The choice of model sampler is dependent on both the cardinality of the model space and the structure of the problem. When $|\pazocal{M}|$ is small, the Gaussian process surrogate-based sampler or the categorical sampler using Monte Carlo gradients are both appropriate, although in practice it is usually safe to default to the latter approach. For high cardinality problems, we recommend a neural model sampler for approximate inference on the distribution of model weights.
@@ Acknowledgments
This research was made possible by funding support from the CSIRO Machine Learning and Artificial Intelligence Future Science Platform. The authors would also like to thank Ryan Thompson for critical advice in the construction of the non-linear DAG example, and Daniel M. Steinberg and Edwin Bonilla for advice on Monte Carlo gradient estimation methods. SAS is supported by the Australian Research Council.